# CQ-DINO: Mitigating Gradient Dilution via Category Queries for Vast Vocabulary Object Detection

### v1

@@ Abstract
With the exponential growth of data, traditional object detection methods are increasingly struggling to handle vast vocabulary object detection tasks effectively. We analyze two key limitations of classification-based detectors: positive gradient dilution, where rare positive categories receive insufficient learning signals, and hard negative gradient dilution, where discriminative gradients are overwhelmed by numerous easy negatives. To address these challenges, we propose CQ-DINO, a category query-based object detection framework that reformulates classification as a contrastive task between object queries and learnable category queries. Our method introduces image-guided query selection, which reduces the negative space by adaptively retrieving top-K relevant categories per image via cross-attention, thereby rebalancing gradient distributions and facilitating implicit hard example mining. Furthermore, CQ-DINO flexibly integrates explicit hierarchical category relationships in structured datasets (e.g., V3Det) or learns implicit category correlations via self-attention in generic datasets (e.g., COCO). Experiments demonstrate that CQ-DINO achieves superior performance on the challenging V3Det benchmark (surpassing previous methods by 2.1% AP) while maintaining competitiveness in COCO. Our work provides a scalable solution for real-world detection systems requiring wide category coverage. The dataset and code will be publicly at https://github.com/RedAIGC/CQ-DINO.
@@
1 Introduction
With the rapid expansion of data, developing a robust AI system capable of large-scale object detection has become essential. This necessity is driven by real-world applications’ increasing complexity and diversity, where AI must manage extensive vocabulary and dynamic environments. Vast vocabularies inherently present hierarchical category structures, as illustrated by classification datasets like ImageNet [7] and Bamboo [48]. Recent detection benchmarks such as V3Det [37], featuring 13,204 object categories organized in hierarchical structures, highlight the magnitude of this challenge. While object detection has witnessed significant advancements [27, 17, 22, 10], scaling effectively to vast vocabularies remains a substantial challenge.
Classification head-based methods typically employ feed-forward networks (FFNs) with sigmoid activation and Focal loss [30] optimization. While these approaches succeed on benchmarks with limited categories such as COCO [19] (80 categories) and Objects365 [31] (365 categories), they face fundamental challenges and scalability in vast vocabulary settings. Recent benchmarks such as V3Det [37], featuring 13,204 object categories, highlight the challenges in vast vocabulary detection. Text-prompted contrastive methods leverage Vision-Language Models (VLMs) to achieve impressive performance in open-vocabulary [45] and visual grounding [44] tasks. These approaches inherit strong generalization capabilities by encoding target categories as text inputs to VLMs. However, they struggle to perform in vast vocabulary scenarios where the text input grows prohibitively long, limiting their practical applicability. Language model-generated methods address open-ended [18] object detection by employing language models to generate category labels without predefined inputs. While theoretically capable of detecting arbitrary objects, these approaches cannot reliably control the granularity of generated categories, creating a fundamental misalignment with practical detection requirements.
@@ 1 Introduction
To tackle these challenges, we introduce category query-based contrastive detection (CQ-DINO), a DINO-based architecture that encodes categories as learnable queries. Our key insight is that dynamic sparse category selection reduces negative space. To this end, we introduce image-guided query selection that identifies relevant categories based on image content. We compute category-to-image similarity through cross-attention layers, followed by top-K selection. This allows us to retain only the most relevant categories for each image, typically a small fraction of the total vocabulary. This selection mechanism provides three crucial benefits: (1) it balances the ratio between positive and negative gradients, (2) it performs implicit hard mining by selecting the most similar categories, and (3) it reduces memory and computational costs, making the framework scalable to extremely large vocabularies. The selected category queries then interact with image features to obtain object queries. From these object queries, bounding boxes are predicted through a cross-modality decoder. Final classifications are obtained using contrastive alignment between object queries and category queries.
The category query offers flexibility beyond traditional classification head-based methods, as it can easily encode inter-category correlations. For datasets with explicit category hierarchies like V3Det [37], we leverage the inherent tree structure to construct hierarchical category queries with an adaptive weight that balances local and hierarchical features. For datasets without explicit hierarchies (e.g., COCO [19]), we use self-attention to learn correlations between categories implicitly.
@@ 2.1 Vast vocabulary object detection
The progression of object detection datasets reveals a trend toward increasingly expansive vocabularies: from small-scale benchmarks like PASCAL VOC [9] (20 classes) to medium-scale datasets such as Objects365 [31] (365 classes), to large-vocabulary benchmarks like LVIS [11] (1,203 classes). Wang et al. [37] introduce the first vast vocabulary object detection dataset, V3Det. V3Det encompasses 13,204 classes organized in hierarchical structures and presents new scalability challenges.
While the LVIS dataset serves as a large-vocabulary detection benchmark, it primarily presents challenges related to long-tail distributions (28% classes appear in fewer than 11 images). To address these challenges, several loss functions have been proposed, including Equalization loss [32, 33, 50] and SeeSaw loss [36], which reweight per-class losses by balancing gradients or sample frequencies. Additionally, Yang et al. [41], and Detic [53] incorporate additional image-level datasets with fine-grained classes to improve detection performance for rare categories.
@@
The V3Det Challenge [38] has inspired several pipelines for vast vocabulary object detection. The MixPLv2 [38] model designs a semi-supervised pipeline that leverages V3Det as the labeled dataset and Objects365 [31] as the unlabeled dataset. RichSem-DINO-FocalNet [38] utilizes the RichSem-DINO detection framework [24], incorporating FocalNet-Huge [42] as its backbone pretrained on the Objects365 dataset [31]. In recent work, Prova [4] proposes specifically for V3Det and introduces detailed image and text prototypes for classification. However, these methods still rely on fundamentally similar classification architectures, raising questions about their viability for even larger vocabularies beyond tens of thousands of categories.
2.2 Classification-based detectors
Object detection architectures typically employ feedforward networks (FFNs) as fixed classification heads for category prediction. Traditional two-stage detectors[27, 2, 12] first generate class-agnostic proposals through Region Proposal Networks (RPNs), then classify these regions using FFN-based classifiers. In contrast, one-stage detectors[30, 35, 8] streamline detection by directly predicting bounding boxes and categories in a single step.
@@ 2.2 Classification-based detectors
The transformer-based DETR [3] reformulated object detection as a set prediction problem using learnable queries. However, it suffered from slow convergence. Subsequent works [54, 39, 20, 16, 46, 25, 21, 13] address these limitations. For example, Deformable DETR [54] proposes multi-scale deformable attention with sparse spatial sampling. DINO [46] achieves impressive performance by leveraging contrastive query denoising and mixed query selection strategies. Despite these architectural advances, existing methods fundamentally rely on classification heads with activation functions, trained using Focal loss [30] or Cross-Entropy loss [49]. This design paradigm inherently limits scalability and poses optimization challenges when scaling up to vast vocabulary scenarios.
@@
With the rapid development of multi-modal models, visual and textual information integration has become increasingly seamless. Recent vision-language models enable text-prompted object detection by aligning visual features with text features. These methods achieve impressive performance in open-vocabulary and visual grounding tasks. GLIP [17] pioneers this approach by establishing contrastive learning between image regions and textual phrases using large-scale image-text pairs. Grounding DINO [22] further enhances cross-modal alignment through the early fusion of vision and language features. Similarly, DetCLIP [43] and RegionCLIP [51] leverage image-text pairs with pseudo-labels to expand region-level semantic understanding for stronger generalization ability.
However, these text-prompted methods face fundamental limitations when scaling to vast vocabularies, primarily due to their restricted text token capacity. For instance, GLIP [17] and Grounding DINO [22] limit input prompts to approximately 128 tokens, which corresponds to roughly 40 categories per inference pass. Consequently, detecting all classes in a dataset like V3Det [37] (13,204 categories) would require over 331 sequential inference passes per image. Therefore, these methods are computationally impractical for real-time applications and are not feasible for large-scale vocabulary object detection.
2.3 Generative detectors
In recent years, the success of generative models has driven significant advancements in multimodal large language models (MLLMs), inspiring detection approaches that leverage their powerful visual understanding and generative capabilities. Although some multi-modal large models (e.g., InternVL [5], Qwen-VL [1], Gemini [34]) have demonstrated preliminary object detection capabilities, they still lack the ability for precise detection. To address this limitation, recent works[18, 10, 15, 37, 40] have utilized LLMs specifically as category generators rather than direct detectors. GenerateU [18] employs a T5 model [6] to decode object categories from visual features, thereby framing detection as a text-generation task. Similarly, LLMDet [10] and ChatRex [15] utilize instruction-tuned LLMs to predict object categories from image features.
Although open detection is theoretically possible, these generation methods have inconsistencies with the detection requirements due to the uncontrollable granularity of the generated categories. For example, when regions contain the term “Persian cat”, these methods may output a more general term “cat”, introducing ambiguity and reducing the accuracy required for fine-grained detection tasks.
Figure 3: Overview of the CQ-DINO framework for vast vocabulary object detection. Key components include: (1) Learnable category queries enhanced with hierarchical tree construction for semantic relationship modeling; (2) Image-guided query selection that identifies the most relevant category queries; (3) Feature enhancer and cross-modality decoder (adapted from GroundingDINO [22]) that processes object queries with contrastive alignment between object queries and selected category queries.
Figure 4: Illustration of image-guided query selection module.
Figure 5: Hierarchical tree construction for category queries.
@@ 3.1 Challenges in vast vocabulary detection
Existing methods for vast vocabulary object detection with $C$ categories ($C>10^{4}$), particularly those employing sigmoid-based classifiers with Focal loss [30], face fundamental optimization challenges. To systematically analyze these issues, we consider a simplified formulation using Cross-Entropy loss with sigmoid activation, revealing two critical limitations:
1) Positive gradient dilution. In vast vocabulary detection, the gradient signal for positive classes is overwhelmed by the aggregated negative gradients. Let $z_{c}$ denote the logit for class $c$ and $y_{c}\in\{0,1\}$ its ground-truth label. The gradient of the Cross-Entropy loss $\mathcal{L}$ with respect to $z_{c}$ is:
@@
The total negative gradient magnitude grows linearly with category count $C$:
@@
As $C$ increases in vast category scenarios (  e.g., $C>10^{4}$), $\rho\to 0$ causes positive gradients to be suppressed during optimization. This fundamentally hinders the model’s ability to learn from positive examples.
@@
Fig. 2 demonstrates these theoretical challenges. The gradient ratio for the V3Det dataset (13,204 classes) is smaller than for the COCO dataset (80 classes), revealing the inherent difficulty in vast vocabulary object detection. While Focal loss partially mitigates these issues by down-weighting easy negatives, the gradient ratio for V3Det remains around 0.5 compared to approximately 1.0 for COCO, indicating that gradient imbalance persists despite these improvements. We provide a comprehensive experimental analysis of Focal loss performance and limitations in Section 4.5.
@@ 3.2 CQ-DINO
2) Image-guided query selection. Given image features $F_{img}\in\mathbb{R}^{B\times D\times H^{\prime}\times W^{\prime}}$ from the image encoder, we compute the similarity between $F_{img}$ and $Q_{cat}$ through multi-head cross-attention modules. For each image, we select the top-$C^{\prime}$ most relevant queries ($C^{\prime}\ll C$, typically $C^{\prime}=100$ for $C>10^{4}$), ensuring that the target class $c^{+}$ and its semantically confusing negative classes are preserved. This selection process rebalances the gradients and implicitly performs hard negative mining.
3) Feature enhancer and cross-modality decoder. The selected category queries $Q_{cat}^{\prime}\in\mathbb{R}^{B\times C^{\prime}\times D}$ along with image features $F_{img}$ are processed through GroundingDINO [22] components. First, the feature enhancer module fuses category queries and image features. Next, object queries are generated through language-guided query selection. Finally, detection outputs are produced by the cross-modality decoder and contrastive alignment between object queries and selected category queries.
@@ 3.3 Image-guided query selection
The core innovation of CQ-DINO is our image-guided category selection module, illustrated in Fig. 4. Our selection module employs cross-attention between category queries $Q_{cat}$ and image features $F_{img}$. Here, $Q_{cat}$ serve as queries (Q), while $F_{img}$ provide keys (K) and values (V). The cross-attention layer establishes category-to-image correlations through similarity computation. Then, we apply TopK selection to retain only the top-$C^{\prime}$ categories ($C^{\prime}\ll C$) based on activation values. We supervise this selection using Asymmetric loss [29], which serves as a multi-class classification loss.
@@ 3.4 Encoding category correlations
A key advantage of our category query approach is the capacity to model complex semantic relationships among categories, which is difficult for traditional classification head-based methods. We propose two complementary strategies for encoding category correlations.
Explicit hierarchical tree construction. For hierarchical datasets like V3Det [37], we introduce hierarchical tree construction, as shown in Fig. 5. The process begins at leaf nodes and progressively integrates hierarchical information upward through the category tree. For each parent node $v$ with children $\mathcal{C}(v)$, its new query $Q_{v}^{\prime}$ is a combination of its original query $Q_{v}$ and the mean pooling of all direct child nodes. Leaf nodes retain their original features directly since they have no children.
@@
where $n_{v}$ is the child count for node $v$, $N_{\text{max}}$ is the maximum child count across the tree, and $w\in[0,0.5]$ is a hyperparameter (default: 0.3). This adaptive weight $\alpha_{v}$ ensures parent nodes with more descendants incorporate more collective knowledge, while nodes with fewer children maintain stronger individual semantics.
Building upon this structural design, we introduce a masking strategy during classification loss computation to mitigate hierarchical ambiguity. If any child category exists in the ground truth, its parent nodes are excluded from the classification loss. This prevents conflicting supervision signals for semantically related categories, such as suppressing “vehicle” when “car” is annotated.
Implicit relation learning. For categories without explicit hierarchical structure, we employ a self-attention mechanism to learn category relationships. This allows semantically related categories to influence each other’s representations based on learned attention patterns.

### v3

@@ Abstract
With the exponential growth of data, traditional object detection methods are increasingly struggling to handle vast vocabulary object detection tasks effectively. We analyze two key limitations of classification-based detectors: positive gradient dilution, where rare positive categories receive insufficient learning signals, and hard negative gradient dilution, where discriminative gradients are overwhelmed by numerous easy negatives. To address these challenges, we propose CQ-DINO, a category query-based object detection framework that reformulates classification as a contrastive task between object queries and learnable category queries. Our method introduces image-guided query selection, which reduces the negative space by adaptively retrieving the top-K relevant categories per image via cross-attention, thereby rebalancing gradient distributions and facilitating implicit hard example mining. Furthermore, CQ-DINO flexibly integrates explicit hierarchical category relationships in structured datasets (e.g., V3Det) or learns implicit category correlations via self-attention in generic datasets (e.g., COCO). Experiments demonstrate that CQ-DINO achieves superior performance on the challenging V3Det benchmark (surpassing previous methods by 2.1% AP) while maintaining competitiveness on COCO. Our work provides a scalable solution for real-world detection systems requiring wide category coverage. The code is available at https://github.com/RedAIGC/CQ-DINO.
1 Introduction
With the rapid expansion of data, developing a robust AI system capable of large-scale object detection has become essential. This necessity is driven by the increasing complexity and diversity of real-world applications, where AI must manage an extensive vocabulary and dynamic environments. Vast vocabularies inherently present hierarchical category structures, as illustrated by classification datasets like ImageNet [7] and Bamboo [43]. Recent detection benchmarks, such as V3Det [35], featuring 13,204 object categories organized in hierarchical structures, highlight the magnitude of this challenge. While object detection has witnessed significant advancements [27, 17, 22, 10], scaling effectively to vast vocabularies remains a substantial challenge.
Category prediction mechanisms can be broadly categorized into four types, as illustrated in Fig. 1. Classification head-based methods employ feed-forward networks (FFNs) with sigmoid activation and Focal Loss [30] optimization. These approaches perform well on benchmarks with limited categories such as COCO [19] (80 categories) and Objects365 [31] (365 categories), but face fundamental challenges and scalability issues in vast vocabulary settings. Text-prompted contrastive methods leverage Vision-Language Models (VLMs) to encode target categories as text inputs, achieving strong open-vocabulary detection [40] performance. However, in vast vocabulary scenarios, text input sequences become prohibitively long, necessitating the splitting of category lists across multiple inference passes. This substantially limits their practical scalability and efficiency. Language model-generated methods approach open-ended [18] object detection by using language models to generate category labels without predefined candidate sets. Although these methods theoretically enable the detection of arbitrary object categories, they generally lack mechanisms to control the granularity of generated labels, which can result in substantial misalignment with practical detection requirements.
@@ 1 Introduction
To tackle these challenges, we introduce Category Query-based DINO (CQ-DINO), a novel architecture that encodes categories as learnable query embeddings. Our approach centers on image-guided query selection, which identifies relevant categories via category-to-image similarity through cross-attention. The key insight driving our method is that dynamic sparse category selection significantly reduces the negative search space. This selection mechanism provides three crucial benefits: (1) balancing the ratio between positive and negative gradients, (2) performing implicit hard mining by selecting the most similar categories, and (3) reducing memory and computational costs, making the framework scalable to extremely large vocabularies. Selected category queries interact with image features to generate object queries. From these object queries, bounding boxes are predicted through a cross-modality decoder. The final classifications are obtained using contrastive alignment between the object queries and the category queries. Unlike traditional classification head-based methods, our category query representation offers greater flexibility by naturally encoding inter-category relationships. For structured datasets with explicit category hierarchies like V3Det [35], we leverage the inherent tree structure to construct hierarchical category queries with an adaptive weighting mechanism that balances local and hierarchical features. For datasets without explicit hierarchies (e.g., COCO [19]), we employ self-attention mechanisms to implicitly learn the correlations between categories.
@@ 2.1 Vast Vocabulary Object Detection
The progression of object detection benchmarks reflects a steady growth in category vocabulary, evolving from relatively small vocabulary datasets such as PASCAL VOC [9] (20 classes) and COCO [19] (80 classes), to larger vocabulary benchmarks, including Objects365 [31] (365 classes) and Open Images [15] (600 classes). Recently, Wang et al. [35] introduce V3Det, the first vast vocabulary object detection dataset, comprising 13,204 hierarchically structured categories. This unprecedented scale poses significant challenges in terms of scalability and representation.
@@
Recent progress has been driven by the V3Det Challenge [34], yielding several methodological advancements. MixPLv2 proposes a semi-supervised framework that combines labeled V3Det data with unlabeled Objects365 [31] images through pseudo-labeling. RichSem-DINO-FocalNet enhances detection robustness by integrating the RichSem-DINO [24] framework with a FocalNet-Huge backbone [38] pretrained on Objects365. Most recently, Prova [4] introduces multi-modal image-text prototypes specifically optimized for V3Det’s fine-grained classification. However, these methods still rely on fundamentally similar classification architectures, raising questions about their effectiveness and scalability for even larger vocabularies beyond tens of thousands of categories.
2.2 Object Detectors
@@
Classification Head-based Methods. Most object detection frameworks employ feedforward networks (FFNs) as fixed classification heads for category prediction. In two-stage detectors [27, 2, 11], region proposals are generated in a class-agnostic manner by Region Proposal Networks (RPNs), followed by FFN-based region classifiers. By contrast, one-stage detectors [30, 33, 8] predict bounding boxes and categories in a single step. Transformer-based methods, such as DETR [3], reformulate detection as a set prediction problem using learnable queries. Although DETR achieves an elegant end-to-end design, it suffers from slow convergence. Subsequent works [47, 36, 20, 16, 41, 25, 21, 12] mitigate these limitations. For example, Deformable DETR [47] proposes multi-scale deformable attention for sparse spatial sampling. DINO [41] improves performance via contrastive query denoising and mixed query selection. Despite these architectural advances, current methods fundamentally rely on classification heads with activation functions, typically optimized with Focal Loss [30] or Cross-Entropy Loss [44]. This design inherently constrains scalability and presents optimization challenges when extending to vast vocabulary detection scenarios.
@@
Text-prompted Contrastive Methods. Vision-language models have advanced the seamless integration of visual and textual modalities for open-vocabulary object detection. These methods encode target categories as text inputs and align visual and textual representations. For instance, GLIP [17] pioneers the use of contrastive learning between image regions and textual phrases. Grounding DINO [22] further improves cross-modal alignment through early fusion of vision and textual features. Similarly, DetCLIP [39] and RegionCLIP [45] leverage image-text pairs with pseudo-labels to enhance region-level semantic understanding and improve generalization. Despite these advances, text-prompted methods face scalability bottlenecks due to the limited capacity of text token inputs during inference. For example, GLIP [17] and Grounding DINO [22] restrict input prompts to approximately 128 tokens per pass, allowing for about 40 categories simultaneously. Thus, detecting all categories in a vast vocabulary benchmark like V3Det [35] (13,204 classes) would require over 331 sequential inference passes per image. This constraint renders current text-prompted methods computationally inefficient and impractical for real-time or large-scale detection.
Language Model-Generated Methods. Advances in multimodal large language models (MLLMs) have inspired detection methods leveraging their visual understanding and generative abilities. Some MLLMs [5, 1, 32] show preliminary object detection abilities but exhibit limited localization precision and recognition granularity. To mitigate these limitations, recent works [18, 10, 14, 35, 37] use LLMs primarily as category generators rather than direct detectors. For instance, GenerateU [18] employs a T5-based decoder [6] to generate category names from visual features, reframing detection as text generation. Similarly, LLMDet [10] and ChatRex [14] utilize instruction-tuned LLMs to predict object categories from image features. While these generative approaches enable open-ended detection, they often produce inconsistencies due to limited controllability over label granularity. For example, given an image region of a “Persian cat”, the model may generate the generic term “cat”, causing semantic ambiguity and reduced accuracy for fine-grained detection tasks.
@@ 3.1 Challenges in Vast Vocabulary Detection
Existing methods for vast vocabulary object detection with $C$ categories ($C>10^{4}$), particularly those employing sigmoid-based classifiers with Focal Loss [30], face fundamental optimization challenges. We systematically analyze these issues through a simplified formulation using the Cross-Entropy Loss with sigmoid activation, revealing two critical limitations:
1) Positive Gradient Dilution. In vast vocabulary detection, the gradient signal for positive classes is overwhelmed by the aggregated negative gradients. Let $z_{c}$ denote the logit for class $c$ and $y_{c}\in\{0,1\}$ be its ground-truth label. The gradient of the Cross-Entropy Loss $\mathcal{L}$ with respect to $z_{c}$ is:
@@
The total negative gradient magnitude grows linearly with the category count $C$:
@@
During early training stages, $\epsilon$ retains a non-negligible value. Since $C~{}\text{exceeds}~{}10^{4}$ in vast category scenarios, $\rho\to 0$, causing positive gradients to be suppressed relative to the cumulative negative gradients. This fundamentally hinders the model’s ability to learn from positive examples.
@@
Fig. 2 demonstrates these theoretical challenges. The gradient ratio for the V3Det dataset (13,204 classes) is lower than for the COCO dataset (80 classes), revealing the inherent difficulty in vast vocabulary object detection. While Focal Loss partially mitigates these issues by down-weighting easy negatives, the gradient ratio for V3Det remains around 0.5 compared to approximately 1.0 for COCO, indicating that gradient imbalance persists despite these improvements. We provide a comprehensive experimental analysis of Focal Loss performance and limitations in Sec. 4.4.
Figure 3: Overview of the CQ-DINO framework for vast vocabulary object detection. Key components: (1) Learnable category queries enhanced with hierarchical tree construction for semantic relationship modeling; (2) Image-guided query selection that identifies the most relevant category queries; (3) Feature enhancer and cross-modality decoder (adapted from GroundingDINO [22]), processing object queries with contrastive alignment between object and selected category queries.
@@ 3.2 CQ-DINO
2) Image-guided query selection. Given image features $F_{img}\in\mathbb{R}^{B\times D\times H^{\prime}\times W^{\prime}}$ from the image encoder, we compute similarities between $F_{img}$ and $Q_{cat}$ through multi-head cross-attention modules. For each image, we select the top-$C^{\prime}$ most relevant queries ($C^{\prime}\ll C$, typically $C^{\prime}=100$ for $C>10^{4}$), ensuring that the target class $c^{+}$ and its most confusing negative classes are preserved. This selection process rebalances the gradients and implicitly performs hard negative mining.
3) Feature enhancer and cross-modality decoder. The selected category queries $Q_{cat}^{\prime}\in\mathbb{R}^{B\times C^{\prime}\times D}$ and image features $F_{img}$ are processed through GroundingDINO [22] components. First, the feature enhancer module fuses the category queries and image features. Next, object queries are generated through language-guided query selection. Finally, detection outputs are produced by the cross-modality decoder and contrastive alignment between object queries and selected category queries.
Figure 4: Illustration of image-guided query
selection module.
Figure 5: Hierarchical tree construction for
category queries.
@@ 3.3 Image-Guided Query Selection
The core innovation of CQ-DINO is our image-guided category selection module, illustrated in Fig. 5. This module employs cross-attention between category queries $Q_{cat}$ and image features $F_{img}$. Here, $Q_{cat}$ serves as queries (Q), while $F_{img}$ provides keys (K) and values (V). The cross-attention layer establishes category-to-image correlations through similarity computation. Then, we apply TopK selection to retain only the top-$C^{\prime}$ categories ($C^{\prime}\ll C$) based on activation values. We supervise this selection using Asymmetric Loss [29], which serves as a multi-class classification loss.
@@ 3.4 Encoding Category Correlations
A key advantage of our category query approach is the capacity to model complex semantic relationships among categories, which is difficult for traditional classification head-based methods to achieve. We propose two complementary strategies for encoding category correlations.
Explicit Hierarchical Tree Construction. For hierarchical datasets like V3Det [35], we introduce hierarchical tree construction, as shown in Fig. 5. The process begins at leaf nodes and progressively integrates hierarchical information upward through the category tree. For each parent node $v$ with children $\mathcal{C}(v)$, its new query $Q_{v}^{\prime}$ is a combination of its original query $Q_{v}$ and the mean pooling of all direct child nodes. Leaf nodes retain their original features directly, since they have no children.
@@
where $n_{v}$ is the child count for node $v$, $N_{\text{max}}$ is the maximum child count across the tree, and $w\in[0,0.5]$ is a hyperparameter (default: 0.3). This adaptive weight $\alpha_{v}$ ensures that parent nodes with more descendants incorporate more collective knowledge, while nodes with fewer children maintain stronger individual semantics.
Building upon this structural design, we introduce a masking strategy during the classification loss computation to mitigate hierarchical ambiguity. If any child category exists in the ground truth, its parent nodes are excluded from the classification loss. This prevents conflicting supervision signals for semantically related categories, such as suppressing “vehicle” when “car” is annotated.
Implicit Relation Learning. For categories without an explicit hierarchical structure, we employ a self-attention mechanism to learn category relationships. This allows semantically related categories to influence each other’s representations based on learned attention patterns.